# An Embedding Theorem for tractor bundles, and an application in conformal pseudo-Riemannian geometry

### v1

@@ 4.2. Consequences of equivariance for stabilizer and isotropy
For a sum $\sigma=-n_{1}\alpha_{1}-n_{2}\alpha_{2}$ with $n_{1}\leq 1$, excluding $-\alpha_{2}$ and $-2\alpha_{2}$ gives $\Sigma^{\prime}$ with no two elements summing to $\sigma$ and corresponding to a subspace of $\mathfrak{h}/\mathfrak{h}^{0}$ of dimension at least 4. Thus we may assume $\rho(2\beta_{1})\neq n_{1}\alpha_{1}+n_{2}\alpha_{2}$ with $n_{1}\leq 1$, nor, by symmetry, with $n_{2}\leq 1$. The only possibility is thus that $\rho(2\beta_{1})=2\alpha_{1}+2\alpha_{2}$, with an isotropic subspace of dimension 3.
In the case (2), the root spaces $\mathfrak{h}_{-\alpha_{i^{\prime}}-\alpha_{j^{\prime}}}$ have trivial intersection with $\mathfrak{h}^{0}$ for all $j^{\prime}$, and Lemma 4.5 (3) implies that $\mathfrak{h}_{-\alpha_{i^{\prime}}}\cap\mathfrak{h}^{0}=0$ as well. For the collection of roots $\Sigma=\{-\alpha_{i^{\prime}}-\alpha_{k^{\prime}},-\alpha_{i^{\prime}}:1\leq k^{\prime}\leq p^{\prime}\}$ any sum of two elements is obtained from a unique unordered pair, and has the form $-2\alpha_{i^{\prime}}-\alpha^{\prime}$ for $\alpha^{\prime}$ as in the conclusion of (2), or $\alpha^{\prime}=\alpha_{i^{\prime}}$ or $2\alpha_{i^{\prime}}$. If $-\rho(2\beta_{1})$ is not such a sum, then all of the root spaces $\mathfrak{h}_{-\alpha_{i^{\prime}}-\alpha_{k^{\prime}}}$ together with $\mathfrak{h}_{-\alpha_{i^{\prime}}}$ comprise an isotropic subspace, and it has dimension $12p^{\prime}+$ 1. If $\rho(2\beta_{1})=2\alpha_{i^{\prime}}+\alpha^{\prime}$ with $\alpha^{\prime}=\alpha_{i^{\prime}}$ or $2\alpha_{i^{\prime}}$, then $\mathfrak{h}_{-2\alpha_{i^{\prime}}}$ can be omitted, leaving an isotropic subspace of dimension $2p^{\prime}$. Thus the maximal isotropic subspace can have dimension $12p^{\prime}-$ 1 only if $\alpha^{\prime}$ is as in the conclusion of (2). ∎
@@
If $\beta_{j}$ is a short root in $\check{\Delta}(2\alpha_{i^{\prime}})$, then it was obtained as the sum of $\beta_{j}-\beta_{i}$ and $\beta_{i}$ in $\check{\Delta}(\alpha_{i^{\prime}})$ for some $i>j$; this is the only way to obtain $\beta_{j}$ as a sum of two positive roots. There are $X,Y\in\mathfrak{h}_{\alpha_{i^{\prime}}}$ with $[X,Y]=Z\neq 0$ and such that
@@
To ensure the first equality we are using that $\mathfrak{g}_{\beta_{j}-\beta_{i}}$ is 1-dimensional, and replacing $Y$ by a linear combination with $X$ if necessary. The last equality relies on (11). Now $\iota([Y,Z])=[\iota(Z),\iota(Y)]=0$ but $[\iota(Z)_{\beta_{j}},\iota(Y)_{\beta_{i}}]\neq 0$. Since $\rho(\beta_{i})=\alpha_{i^{\prime}}$ and $\rho(\beta_{j})=2\alpha_{i^{\prime}}$, the components $\iota(Y)_{\beta_{j}}$ and $\iota(Z)_{\beta_{i}}$ are both 00. To obtain $\beta_{i}+\beta_{j}$ as the sum of two positive roots of $\Delta_{G}$, these could be $\beta_{i}$ and $\beta_{j}$, or an initial plus a terminal root, of the form $(\beta_{i}-\beta_{k})+(\beta_{j}+\beta_{k})$ for some $k>i$, or vice-versa. In order that the nonzero bracket $[\iota(Z)_{\beta_{j}},\iota(Y)_{\beta_{i}}]$ cancels, there must be a nonzero component of $\iota(Z)\in\check{\mathfrak{h}}_{2\alpha_{i^{\prime}}}$ on one of these initial or terminal roots, yielding an initial or terminal root in $\check{\Delta}(2\alpha_{i^{\prime}})$, as claimed. ∎
@@
The bound $p^{\prime}=1$ follows from Zimmer [40] and Bader–Nevo [4, Thm 2]; see also [28]. It is shown on [29, p. 9] that ${\mathfrak{b}}_{H}\subset\mathfrak{h}^{0}$.
@@
First, suppose $\mathfrak{h}_{-\alpha_{i^{\prime}}}$ is contained in $\mathfrak{h}^{0}$ for all $i^{\prime}$. The ad $\mbox{ad }{\mathfrak{b}}_{H}$-invariance of $\mathfrak{h}^{0}$ then implies that all nonnegative root spaces are in $\mathfrak{h}^{0}$, as well. Now ${\mathfrak{b}}_{H}$ and all $\mathfrak{h}_{\pm\alpha_{i^{\prime}}}$ are in $\mathfrak{h}^{0}$. These subspaces generate $\mathfrak{h}$, so $\mathfrak{h}=\mathfrak{h}^{0}$. Lemma 4.4 (1) implies that by conjugating in $P$, we may assume $\iota(\mathfrak{h})\subset{\mathfrak{q}}_{ss}\cong\mathfrak{so}(2,q)$, see Remark 3.2. The conclusions of the case $p=1$ apply—note that now $\mathfrak{h}$ acts conformally on the Lorentzian Möbius space of dimension $q$. Thus $p^{\prime}=1$.
Next suppose that one $\overline{\mathfrak{h}}_{-\alpha_{i^{\prime}}}\neq 0$ but that $\mathfrak{h}_{-\alpha_{j^{\prime}}}\subset\mathfrak{h}^{0}$ for all $j^{\prime}\neq i^{\prime}$. An isotropic subspace for $[s]$ has dimension at most $2$; on the other hand, Lemma 4.6 (2), gives the lower bound $12p^{\prime}-$ 1 if $p^{\prime}>1$, so $p^{\prime}=1$.
@@
Suppose that both $\overline{\mathfrak{h}}_{-\alpha_{1}}$ and $\overline{\mathfrak{h}}_{-\alpha_{2}}$ are nonzero. An $[s]$-isotropic subspace has dimension $\leq 3$, so Lemma 4.6 (1) implies $\rho(\beta_{1})=\alpha_{1}+\alpha_{2}$. Then a maximal isotropic subspace for $[s]$ is $\overline{\mathfrak{h}}_{-2\alpha_{1}}+\overline{\mathfrak{h}}_{-\alpha_{1}}+\overline{\mathfrak{h}}_{-\alpha_{2}}$, and all three summands are 1-dimensional. Then $\mathfrak{h}_{-\alpha_{k^{\prime}}}\cap\mathfrak{h}^{0}\neq 0$ for $k^{\prime}=1,2$. From the ad $\mbox{ad }{\mathfrak{b}}_{H}$-invariance of $\mathfrak{h}^{0}$, we obtain ${\mathfrak{a}}_{H}\subset\mathfrak{h}^{0}$ and therefore $\mathfrak{h}_{\alpha_{k^{\prime}}}\subset\mathfrak{h}^{0}$ for $k^{\prime}=1,2$ by Lemma 4.5 (1).
Now by Lemma 4.8 (2), $\check{\Delta}(2\alpha_{k^{\prime}})$ contains an initial or terminal root for both $k^{\prime}=1,2$. They cannot both contain only initial roots by Lemma 4.8 (1) because for any string of three initial roots in $\Delta^{+}_{G}$, there is no further string of three initial roots in its complement. On the other hand $\rho(\beta_{1})=\alpha_{1}+\alpha_{2}$ means that $\check{\Delta}(\alpha_{1})$ and $\check{\Delta}(\alpha_{2})$ do not both contain only short roots. Therefore, for $\ell,j,m>1$ all distinct,
@@
$\{\beta_{\ell},\beta_{j}\}\subseteq\check{\Delta}(\alpha_{k^{\prime}})\subseteq\rho^{-1}(\alpha_{k^{\prime}})$$\{\beta_{\ell}+\beta_{j}\}\subseteq\check{\Delta}(2\alpha_{k^{\prime}})\subseteq\rho^{-1}(2\alpha_{k^{\prime}})$
$\{\beta_{1}-\beta_{\ell},\beta_{\ell}-\beta_{m}\}\subseteq\check{\Delta}(\alpha_{\ell^{\prime}})\subseteq\rho^{-1}(\alpha_{\ell^{\prime}})$$\{\beta_{1}-\beta_{m}\}\subseteq\check{\Delta}(2\alpha_{\ell^{\prime}})\subseteq\rho^{-1}(2\alpha_{\ell^{\prime}})$
These values determine $\rho$, which means that all the above containments are in fact equalities. By Lemma 4.7, $\rho^{-1}(\alpha_{1}-\alpha_{2})$ must contain at least one positive root; this is only possible if $k^{\prime}=1$ and $\ell^{\prime}=2$. Also $\check{\Delta}(\alpha_{1}+\alpha_{2})=\{\beta_{1}\}$.
Let $X\in\mathfrak{h}_{\alpha_{2}}$ and $Y\in\mathfrak{h}_{\alpha_{1}}$ with $[X,Y]=Z\neq 0$. From equation (11), $[\iota(Y),\iota(X)]=\iota(Z)$. Now $\iota(Z)\in\mathfrak{g}_{\beta_{1}}$, and it equals $[\iota(Y)_{\beta_{\ell}},\iota(X)_{\beta_{1}-\beta_{\ell}}]$. But also by (11) and the properties of strings recounted above,
$\iota\left([Z,Y]\right)=[\iota(Y),\iota(Z)]=[\iota(Y)_{\beta_{\ell}}+\iota(Y)_{\beta_{j}},\iota(Z)_{\beta_{1}}]\neq 0$
because the component on $\beta_{\ell}+\beta_{1}$ is nonzero. But $[Z,Y]=0$, a contradiction. ∎
@@
Now Lemma 4.6 implies $p\geq 2p^{\prime}-1$. Suppose they are equal. If $p=3$, then $p^{\prime}=2$, and Proposition 4.11 gives that $\mathfrak{h}_{-\alpha_{j^{\prime}}}\subset\mathfrak{h}^{0}$ for all $j^{\prime}\neq i^{\prime}$; otherwise, for $p>3$, Lemma 4.6 (1) gives the same conclusion. The ad $\mbox{ad }{\mathfrak{b}}_{H}$-invariance of $\mathfrak{h}^{0}$ from Proposition 4.3 (1) implies that $\mathfrak{h}_{\alpha_{j^{\prime}}}\subset\mathfrak{h}^{0}$ for all $j^{\prime}\neq i^{\prime}$. By Lemma 4.5 (3), $\mathfrak{h}_{-\alpha_{i^{\prime}}}\cap\mathfrak{h}^{0}=0$. Now let $\rho(2\beta_{1})=2\alpha_{i^{\prime}}+\alpha^{\prime}$ as in Lemma 4.6 (2).
Case 1: $\rho(\beta_{1})=\alpha_{i^{\prime}}$. As was seen in the proof of Lemma 4.6, root spaces are orthogonal unless the roots add up to the conformal factor, in this case $-2\alpha_{i^{\prime}}$. Therefore the subspace $\oplus_{j^{\prime}}{\mathfrak{h}}_{-\alpha_{i^{\prime}}-\alpha_{j^{\prime}}}$ projects modulo $\mathfrak{h}^{0}$ to an isotropic subspace for $[s]$; by Lemma 4.5 (2), it has trivial interesection with $\mathfrak{h}^{0}$, so the projection has dimension $2p^{\prime}-1=p$. Any $\overline{\mathfrak{h}}_{\alpha_{i^{\prime}}-\alpha_{j^{\prime}}}$, with $j^{\prime}\neq i^{\prime}$, could be added to make a larger isotropic subspace; they must therefore be 00—that is $\mathfrak{h}_{\alpha_{i^{\prime}}-\alpha_{j^{\prime}}}\subset\mathfrak{h}^{0}$ for all $j^{\prime}\neq i^{\prime}$. Because $[\mathfrak{h}_{\alpha_{j^{\prime}}},\mathfrak{h}_{\alpha_{i^{\prime}}-\alpha_{j^{\prime}}}]=\mathfrak{h}_{\alpha_{i^{\prime}}}$, Proposition 4.3 (1) implies $\mathfrak{h}_{\alpha_{i^{\prime}}}\subset\mathfrak{h}^{0}$ as well.
@@
Because $\iota$ intertwines ad $\mbox{ad }{\mathfrak{a}}_{H}$ with ad $\mbox{ad }\check{{\mathfrak{a}}}$, the image $\iota(J)$ is a sum of subspaces of weight spaces for $\check{{\mathfrak{a}}}$, each intersecting ${\mathfrak{p}}$ trivially, therefore contained in $\mathfrak{g}_{-}$. The image $\iota(I)$ is a maximal isotropic subspace of $\mathfrak{g}_{-}\cong{\bf R}^{p,q}$. The weights for ${\mathfrak{a}}_{G}$ on $\mathfrak{g}_{-}$ are
$\Sigma=\{-\beta_{1}\pm\beta\ :\ \beta=0,\beta_{2},\ldots,\beta_{p+1}\}$
Let $\Sigma_{I}$ comprise the roots in $\Sigma$ for which $\iota(I)$ has nontrivial projection on the corresponding root space. Elements of $\Sigma_{I}$ must have the additional property that $\rho(-\beta_{1}\pm\beta)=-\alpha_{i^{\prime}}-\alpha_{j^{\prime}}$ for some $j^{\prime}$; therefore $-\beta_{1}\notin\Sigma_{I}$. A maximal isotropic subspace of $\oplus_{i>1}\mathfrak{g}_{-\beta_{1}\pm\beta_{i}}$ is of the form

### v2

@@ 4.2. Consequences of equivariance for stabilizer and isotropy
For a sum $\sigma=-n_{1}\alpha_{1}-n_{2}\alpha_{2}$ with $n_{1}\leq 1$, excluding $\alpha_{2}$ and $2\alpha_{2}$ gives $\Sigma^{\prime}$ with no two elements summing to $\sigma$ and corresponding to a subspace of $\mathfrak{h}/\mathfrak{h}^{0}$ of dimension at least 4. Thus we may assume $\rho(2\beta_{1})\neq n_{1}\alpha_{1}+n_{2}\alpha_{2}$ with $n_{1}\leq 1$, nor, by symmetry, with $n_{2}\leq 1$. The only possibility is thus that $\rho(2\beta_{1})=2\alpha_{1}+2\alpha_{2}$, with an isotropic subspace of dimension 3.
In the case (2), the root spaces $\mathfrak{h}_{-\alpha_{i^{\prime}}-\alpha_{j^{\prime}}}$ have trivial intersection with $\mathfrak{h}^{0}$ for all $j^{\prime}$, and Lemma 4.5 (3) implies that $\mathfrak{h}_{-\alpha_{i^{\prime}}}\cap\mathfrak{h}^{0}=0$ as well. For the collection of roots $\Sigma=\{-\alpha_{i^{\prime}}-\alpha_{k^{\prime}},-\alpha_{i^{\prime}}:1\leq k^{\prime}\leq p^{\prime}\}$ any sum of two elements is obtained from a unique unordered pair, and has the form $2\alpha_{i^{\prime}}-\alpha^{\prime}$ for $\alpha^{\prime}$ as in the conclusion of (2), or $\alpha^{\prime}=\alpha_{i^{\prime}}$ or $2\alpha_{i^{\prime}}$. If $\rho(2\beta_{1})$ is not such a sum, then all of the root spaces $\mathfrak{h}_{-\alpha_{i^{\prime}}-\alpha_{k^{\prime}}}$ together with $\mathfrak{h}_{-\alpha_{i^{\prime}}}$ comprise an isotropic subspace, and it has dimension $2p^{\prime}+1$. If $\rho(2\beta_{1})=2\alpha_{i^{\prime}}+\alpha^{\prime}$ with $\alpha^{\prime}=\alpha_{i^{\prime}}$ or $2\alpha_{i^{\prime}}$, then $\mathfrak{h}_{-2\alpha_{i^{\prime}}}$ can be omitted, leaving an isotropic subspace of dimension $2p^{\prime}$. Thus the maximal isotropic subspace can have dimension $2p^{\prime}-1$ only if $\alpha^{\prime}$ is as in the conclusion of (2). ∎
@@
Part (2) is vacuously true if $p=q$, since all roots in $\mathfrak{g}$ are initial or terminal. If $\beta_{j}$ is a short root in $\check{\Delta}(2\alpha_{i^{\prime}})$, then it was obtained as the sum of $\beta_{j}-\beta_{i}$ and $\beta_{i}$ in $\check{\Delta}(\alpha_{i^{\prime}})$ for some $i>j$; this is the only way to obtain $\beta_{j}$ as a sum of two positive roots. There are $X,Y\in\mathfrak{h}_{\alpha_{i^{\prime}}}$ with $X,Y]=Z\neq 0$ and such that
@@
To ensure the first equality we are using that $\mathfrak{g}_{\beta_{j}-\beta_{i}}$ is 1-dimensional, and replacing $Y$ by a linear combination with $X$ if necessary. The last equality relies on (11). Now $\iota([Y,Z])=[\iota(Z),\iota(Y)]=0$ but $\iota(Z)_{\beta_{j}},\iota(Y)_{\beta_{i}}]\neq 0$. Since $\rho(\beta_{i})=\alpha_{i^{\prime}}$ and $\rho(\beta_{j})=2\alpha_{i^{\prime}}$, the components $\iota(Y)_{\beta_{j}}$ and $\iota(Z)_{\beta_{i}}$ are both 0. To obtain $\beta_{i}+\beta_{j}$ as the sum of two positive roots of $\Delta_{G}$, these could be $\beta_{i}$ and $\beta_{j}$, or an initial plus a terminal root, of the form $(\beta_{i}-\beta_{k})+(\beta_{j}+\beta_{k})$ for some $k>i$, or vice-versa. In order that the nonzero bracket $\iota(Z)_{\beta_{j}},\iota(Y)_{\beta_{i}}]$ cancels, there must be a nonzero component of $\iota(Z)\in\check{\mathfrak{h}}_{2\alpha_{i^{\prime}}}$ on one of these initial or terminal roots, yielding an initial or terminal root in $\check{\Delta}(2\alpha_{i^{\prime}})$, as claimed. ∎
@@
Note that $p<q$ in this case because $\mbox{dim }M\geq 3$. The bound $p^{\prime}=1$ follows from Zimmer [40] and Bader–Nevo [4, Thm 2]; see also [28]. It is shown on [29, p. 9] that ${\mathfrak{b}}_{H}\subset\mathfrak{h}^{0}$.
@@
First, suppose $\mathfrak{h}_{-\alpha_{i^{\prime}}}$ is contained in $\mathfrak{h}^{0}$ for all $i^{\prime}$. The $\mbox{ad }{\mathfrak{b}}_{H}$-invariance of $\mathfrak{h}^{0}$ then implies that all nonnegative root spaces are in $\mathfrak{h}^{0}$, as well. Now ${\mathfrak{b}}_{H}$ and all $\mathfrak{h}_{\pm\alpha_{i^{\prime}}}$ are in $\mathfrak{h}^{0}$. These subspaces generate $\mathfrak{h}$, so $\mathfrak{h}=\mathfrak{h}^{0}$. Lemma 4.4 (1) implies that by conjugating in $P$, we may assume $\iota(\mathfrak{h})\subset{\mathfrak{q}}_{ss}\cong\mathfrak{so}(2,q)$, see Remark 3.2. Provided $q\geq 3$, the conclusions of the case $p=1$ apply—note that now $\mathfrak{h}$ acts conformally on the Lorentzian Möbius space of dimension $q$. On the other hand, if $p=q=2$, then there is no nontrivial Lie algebra homomorphism from $\mathfrak{su}(p^{\prime},q^{\prime})$ to $\mathfrak{so}(2,2)\cong\mathfrak{so}(1,2)\oplus\mathfrak{so}(1,2)$. Thus $p^{\prime}=1$.
Next suppose that one $\overline{\mathfrak{h}}_{-\alpha_{i^{\prime}}}\neq 0$ but that $\mathfrak{h}_{-\alpha_{j^{\prime}}}\subset\mathfrak{h}^{0}$ for all $j^{\prime}\neq i^{\prime}$. An isotropic subspace for $[s]$ has dimension at most $2$; on the other hand, Lemma 4.6 (2) gives the lower bound $2p^{\prime}-1$ if $p^{\prime}>1$, so $p^{\prime}=1$.
@@
Suppose that both $\overline{\mathfrak{h}}_{-\alpha_{1}}$ and $\overline{\mathfrak{h}}_{-\alpha_{2}}$ are nonzero. By Lemma 4.5(2),
$\mathfrak{h}_{-\alpha_{1}-\alpha_{2}}\cap\mathfrak{h}^{0}=\mathfrak{h}_{-2\alpha_{k^{\prime}}}\cap\mathfrak{h}^{0}=0,\ \ k^{\prime}=1,2$
@@
An $[s]$-isotropic subspace has dimension $\leq 3$, so Lemma 4.6 (1) implies $\rho(\beta_{1})=\alpha_{1}+\alpha_{2}$. A maximal isotropic subspace for $[s]$ is $\overline{\mathfrak{h}}_{-2\alpha_{1}}+\overline{\mathfrak{h}}_{-\alpha_{1}}+\overline{\mathfrak{h}}_{-\alpha_{2}}$, and all three summands are 1-dimensional. Then $\mathfrak{h}_{-\alpha_{k^{\prime}}}\cap\mathfrak{h}^{0}\neq 0$ for $k^{\prime}=1,2$. From the $\mbox{ad }{\mathfrak{b}}_{H}$-invariance of $\mathfrak{h}^{0}$, we obtain ${\mathfrak{a}}_{H}\subset\mathfrak{h}^{0}$ and therefore $\mathfrak{h}_{\alpha}\subset\mathfrak{h}^{0}$ for all $\alpha\in\Delta_{H}^{+}$ by Lemma 4.5 (1). Using Lemma 4.4, we will henceforth assume $\check{\mathfrak{h}}_{\alpha}=\iota(\mathfrak{h}_{\alpha})$ for all $\alpha\in\Delta_{H}^{+}$.
For $k^{\prime}=1,2$, let
$I_{k^{\prime}}=\bigoplus_{i^{\prime}=1,2}\mathfrak{h}_{-\alpha_{i^{\prime}}}\oplus\mathfrak{h}_{-2\alpha_{k^{\prime}}}$
The projections $\overline{\iota(I_{k^{\prime}})}$ are maximal isotropic subspaces of $\mathfrak{g}/{\mathfrak{p}}$ for $k^{\prime}=1,2$, but they are not mutually orthogonal. The weights for ${\mathfrak{a}}_{G}$ on $\mathfrak{g}/{\mathfrak{p}}\cong{\bf R}^{p,q}$ are
$\Sigma=\{-\beta_{1}\pm\beta\ :\ \beta=0,\beta_{2},\beta_{3},\beta_{4}\};$
if $p=q$ then $\beta=0$ does not occur. The projected root spaces $\overline{\mathfrak{h}}_{\alpha}$ for ${\mathfrak{a}}_{H}$ have images $\overline{\iota(\mathfrak{h}_{\alpha})}$ which decompose into the sum of their intersections with the weight spaces of $\check{{\mathfrak{a}}}$ in $\mathfrak{g}/{\mathfrak{p}}$. Since $\rho(-\beta_{1})=-\alpha_{1}-\alpha_{2}$ and $\overline{\mathfrak{h}}_{\alpha_{k^{\prime}}}$ is $1$-dimensional for $k^{\prime}=1,2$, there are $i,j>1$ such that
$\overline{\iota(\mathfrak{h}_{-\alpha_{1}})}\subset\overline{\mathfrak{g}}_{-\beta_{1}\pm\beta_{i}}\qquad\mbox{and}\qquad\overline{\iota(\mathfrak{h}_{-\alpha_{2}})}\subset\overline{\mathfrak{g}}_{-\beta_{1}\pm\beta_{j}}$
Therefore $\rho(\beta_{i})=\pm\alpha_{2}$ and $\rho(\beta_{j})=\pm\alpha_{1}$; in particular $i\neq j$.
The images $\overline{\iota(\mathfrak{h}_{-2\alpha_{k^{\prime}}})}$ do not meet $\overline{\mathfrak{g}}_{-\beta_{1}}$ for $k^{\prime}=1,2$, also because $\rho(-\beta_{1})=-\alpha_{1}-\alpha_{2}$. A maximal isotropic subspace of $\oplus_{i>1}\overline{\mathfrak{g}}_{-\beta_{1}\pm\beta_{i}}$ is of the form
(12) $\left(\bigoplus_{k\in A}\overline{\mathfrak{g}}_{-\beta_{1}+\beta_{k}}\right)\oplus\left(\bigoplus_{\ell\in B}\overline{\mathfrak{g}}_{-\beta_{1}-\beta_{\ell}}\right)$
for $A\sqcup B=\{2,3,4\}$, since the scalar product gives a nondegenerate pairing of each $\overline{\mathfrak{g}}_{-\beta_{1}-\beta_{k}}$ with $\overline{\mathfrak{g}}_{-\beta_{1}+\beta_{k}}$. There is therefore $k\neq 1,i,j$ such that
$\overline{\iota(\mathfrak{h}_{-2\alpha_{k^{\prime}}})}\subset\overline{\mathfrak{g}}_{-\beta_{1}-\beta_{k}}\qquad\mbox{and}\qquad\overline{\iota(\mathfrak{h}_{-2\alpha_{\ell^{\prime}}})}\subset\overline{\mathfrak{g}}_{-\beta_{1}+\beta_{k}}\qquad\{k^{\prime},\ell^{\prime}\}=\{1,2\}$
Now all values of $\rho$ are determined, with $\rho(\beta_{k})=\alpha_{k^{\prime}}-\alpha_{\ell^{\prime}}$. Note that because $\overline{\mathfrak{h}}_{-\alpha_{1}-\alpha_{2}}\neq 0$ and $\beta_{1}$ is the only possible weight in $\Sigma$ mapping to $\alpha_{1}-\alpha_{2}$, it must indeed be a weight, and $p<q$.
By Lemma 4.8 (2), each $\check{\Delta}(2\alpha_{m^{\prime}})$, for $m^{\prime}=1,2$, contains an initial or terminal root. From the determined values of $\rho$,
$\beta_{1}+\beta_{k}\in\check{\Delta}(2\alpha_{k^{\prime}})\qquad\beta_{1}-\beta_{k}\in\check{\Delta}(2\alpha_{\ell^{\prime}})$
By Lemma 4.8 (1), each of these roots is obtained as a sum of two distinct roots of $\check{\Delta}(\alpha_{m^{\prime}})$ for $m^{\prime}=1,2$. In particular,
$\beta_{1}-\beta_{k}=(\beta_{1}-\beta_{i})+(\beta_{i}-\beta_{k})=(\beta_{1}-\beta_{j})+(\beta_{j}-\beta_{k})$
are the two possibilities for the summands from $\check{\Delta}(\alpha_{\ell^{\prime}})$. The value $\rho(\beta_{1}-\beta_{i})=\alpha_{\ell^{\prime}}$ only if $\rho(\beta_{i})=\alpha_{2}$ and $\ell^{\prime}=1$; then also $\rho(\beta_{j}-\beta_{k})=\alpha_{1}$. The alternative $\ell^{\prime}=2$ is equally possible provided $\rho(\beta_{j})=\alpha_{1}$. We will assume $\ell^{\prime}=1$ for the remainder of the proof, which is easily adapted for $\ell^{\prime}=2$.
Next, there are two possible decompositions into summands from $\check{\Delta}(\alpha_{2})$:
$\beta_{1}+\beta_{k}=(\beta_{1})+(\beta_{k})=(\beta_{1}-\beta_{j})+(\beta_{j}+\beta_{k})$
@@
but since $\rho(\beta_{1}),\rho(\beta_{k})\neq\alpha_{2}$, the only possibility is $\beta_{1}-\beta_{j},\beta_{j}+\beta_{k}\in\check{\Delta}(\alpha_{2})$. Thus $\rho(\beta_{j})=\alpha_{1}$.
Observe that $\check{\Delta}(2\alpha_{2})=\{\beta_{1}+\beta_{k}\}$, as there are no other elements of $\Delta_{G}^{+}$ in $\rho^{-1}(2\alpha_{2})$. Let $X\in\mathfrak{h}_{2\alpha_{2}}$, so $\iota(X)\in\mathfrak{g}_{\beta_{1}+\beta_{k}}$. Let $Y\in\mathfrak{h}_{\alpha_{1}}$ with $\iota(Y)_{\beta_{i}-\beta_{k}}\neq 0$. Then
$\iota(X),\iota(Y)]_{\beta_{1}+\beta_{i}}=[\iota(X)_{\beta_{1}+\beta_{k}},\iota(Y)_{\beta_{i}-\beta_{k}}]\neq 0$
Therefore $\iota(X),\iota(Y)]\neq 0$, which implies by (11) and injectivity of $\iota$ that $X,Y]\neq 0$. But this is a contradiction because $2\alpha_{2}+\alpha_{1}\notin\Delta_{H}$. ∎
@@
Lemma 4.6 implies $p\geq 2p^{\prime}-1$. Suppose they are equal. If $p=3$, then $p^{\prime}=2$, and Proposition 4.11 gives that $\mathfrak{h}_{-\alpha_{j^{\prime}}}\subset\mathfrak{h}^{0}$ for all $j^{\prime}\neq i^{\prime}$; otherwise, for $p>3$, Lemma 4.6 (1) gives the same conclusion. The $\mbox{ad }{\mathfrak{b}}_{H}$-invariance of $\mathfrak{h}^{0}$ from Proposition 4.3 (1) implies that $\mathfrak{h}_{\alpha_{j^{\prime}}}\subset\mathfrak{h}^{0}$ for all $j^{\prime}\neq i^{\prime}$. By Lemma 4.5 (3), $\mathfrak{h}_{-\alpha_{i^{\prime}}}\cap\mathfrak{h}^{0}=0$. Now let $\rho(2\beta_{1})=2\alpha_{i^{\prime}}+\alpha^{\prime}$ as in Lemma 4.6 (2).
Case 1: $\rho(\beta_{1})=\alpha_{i^{\prime}}$. As seen above, root spaces are $[s]$-orthogonal unless the roots add up to the conformal factor, in this case $2\alpha_{i^{\prime}}$. Therefore the subspace $\oplus_{j^{\prime}}{\mathfrak{h}}_{-\alpha_{i^{\prime}}-\alpha_{j^{\prime}}}$ projects modulo $\mathfrak{h}^{0}$ to an isotropic subspace for $[s]$; by Lemma 4.5 (2), it has trivial intersection with $\mathfrak{h}^{0}$, so the projection has dimension $2p^{\prime}-1=p$. Any $\overline{\mathfrak{h}}_{\alpha_{i^{\prime}}-\alpha_{j^{\prime}}}$, with $j^{\prime}\neq i^{\prime}$, could be added to make a larger isotropic subspace; they must therefore be 0—that is $\mathfrak{h}_{\alpha_{i^{\prime}}-\alpha_{j^{\prime}}}\subset\mathfrak{h}^{0}$ for all $j^{\prime}\neq i^{\prime}$. Because $\mathfrak{h}_{\alpha_{j^{\prime}}},\mathfrak{h}_{\alpha_{i^{\prime}}-\alpha_{j^{\prime}}}]=\mathfrak{h}_{\alpha_{i^{\prime}}}$, Proposition 4.3 (1) implies $\mathfrak{h}_{\alpha_{i^{\prime}}}\subset\mathfrak{h}^{0}$ as well.
@@
As above, the image $\overline{\iota(J)}$ is a sum of $\check{{\mathfrak{a}}}$-weight spaces in $\mathfrak{g}/{\mathfrak{p}}$, and the image $\overline{\iota(I)}$ is a maximal isotropic subspace of $\mathfrak{g}/{\mathfrak{p}}\cong{\bf R}^{p,q}$. The weights for ${\mathfrak{a}}_{G}$ on $\mathfrak{g}/{\mathfrak{p}}$, for $p<q$, are
$\Sigma=\{-\beta_{1}\pm\beta\ :\ \beta=0,\beta_{2},\ldots,\beta_{p+1}\},$
and when $p=q$ then $\beta=0$ does not occur. Let $\Sigma_{I}$ comprise the roots in $\Sigma$ for which $\overline{\iota(I)}$ has nontrivial projection on the corresponding ${\mathfrak{a}}_{G}$-weight space. Elements of $\Sigma_{I}$ must have the additional property that $\rho(-\beta_{1}\pm\beta)=-\alpha_{i^{\prime}}-\alpha_{j^{\prime}}$ for some $j^{\prime}$; therefore $\beta_{1}\notin\Sigma_{I}$. As in the previous proof, a maximal isotropic subspace of $\oplus_{i>1}\mathfrak{g}_{-\beta_{1}\pm\beta_{i}}$ is of the form